# MVNN: A Measure-Valued Neural Network for Learning McKean-Vlasov Dynamics from Particle Data

### 01 Apr 2026 v1

@@ 1 Introduction
Interacting particle and agent systems are widely used to model collective dynamics in physics, biology, and the social sciences [10, 1, 35, 74, 69]. Since first-principles approaches to modeling interacting forces are often challenging, a key problem is to infer the underlying interaction mechanisms from trajectory observations. Recent data-driven methods address this by assuming a binary-interaction ansatz and estimating a pairwise interaction kernel via least-squares fitting of observed (finite-difference) velocities as a function of pairwise distances [47, 43, 37, 26, 19, 51]. However, the pairwise assumption is often insufficient for complex, realistic systems. In many biological, social, and physical scenarios, the effective dynamics are governed by a drift of the mean field form, a nonlinear functional of the population distribution, rather than a superposition of two-body forces. Popular examples include crowd dynamics governed by local density constraints rather than individual repulsion [52], vehicular traffic where speed is a nonlinear functional of the flow density [40], and cell migration influenced by chemical concentration fields [33]. Restricting the hypothesis class to pairwise interaction kernels can therefore limit the ability to capture emergent collective behaviors. Beyond modeling fidelity, computational efficiency is another major difficulty since direct simulation of interacting particle systems typically scales as $O(N^{2})$ in the number of particles. Mean-field limits and related numerical techniques, such as the random batch method [30, 23], provide scalable approximations for large populations [32, 25, 70] and have been used in aggregation and swarming, Motsch-Tadmor dynamics, and optimization [6, 13, 12, 2, 58, 11, 14, 29, 49]. This motivates the development of drift models whose evaluation scales linearly with $N$, by utilizing the symmetry arising from agent indistinguishability.
@@
where $f_{t}=\mathcal{L}(\mathbf{X}_{t})$ describes the law of the $\mathbf{X}_{t}$ and $\mathbf{B}_{t}$ is a $d\text{-}$dimensional Wiener process. Numerically, the McKean–Vlasov dynamics can be approximated by a system of interacting particles $(\mathbf{X}_{t}^{i,N})_{1\leq i\leq N}$ with a sufficiently large number of particles $N$. Formally, the initial condition $(\mathbf{X}_{0}^{i,N})_{1\leq i\leq N}$ are independent and identically distributed with law $\mu_{0}$ and each particle evolves according to:
@@ 2 Learning McKean–Vlasov Drifts from Particle Data
We consider a standard interacting particle/agent system as a motivating example. Note that the learning framework developed here does not rely on the pairwise interaction kernel and can be applied to more complex interactions. Consider $N$ agents with states $\mathbf{X}_{t}^{i}\in\mathbb{R}^{d}$ and the (possibly stochastic) dynamics:
@@ 2.1 Measure-Valued Neural Network
for some test functions $g_{i}$ and a finite-dimensional map $f$ (e.g. polynomial/smooth), see, e.g., [27, 61]. On compact subsets of Wasserstein space, such cylindrical classes are dense in appropriate smoothness spaces; see Lemma 3.12 and Definition 2.4 in [27] for a representative result. The key computational advantage is that $\langle g,\mu_{t}^{N}\rangle$ can be evaluated by averaging over particles, which is permutation invariant and scales linearly in $N$. Motivated by this formulation, we approximate the drift by a composition of two neural networks: (1) an embedding network and (2) an interaction network. We define the MVNN drift as
@@
The model jointly learns (i) a finite collection of test functions $\varphi_{\mathrm{emb}}$ that extract informative features from the measure, and (ii) an interaction map $\varphi_{\mathrm{int}}$ that aggregates these features with the local state to approximate the drift. The embedding part can be interpreted as learning an optimal, data-driven test function basis on the space of measures. The embedding network $\varphi_{\mathrm{emb}}$ plays the role of the functions $g_{i}$ in the cylindrical functional representation, but unlike polynomial or finite-element bases, it is learned from data instead of fixed a priori. Given the MVNN drift (2), the associated particle system with $N$ agents is governed by:
@@
Although Proposition 2.1 is a standard consequence of Lipschitz continuity for McKean-Vlasov SDEs, it guarantees that our learned MVNN drift induces a well-posed nonlinear Fokker-Planck evolution at the continuum level. This establishes our method as a rigorous framework for learning continuous governing equations directly from discretized particle-level observations, without requiring any smoothing or filtering procedures, in contrast to existing approaches [55, 21]. However, to complete the theoretical justification, we must demonstrate that the learned $N$-particle system (4) converges to this identified mean-field model (5) as $N\to\infty$.
@@
Let the assumptions of Proposition 2.1 hold. Let $(\mathbf{X}_{t}^{\theta,i,N})_{i=1}^{N}$ be the $N$-particle system solving (4) with $f_{0}$-chaotic initial data $\mathbf{X}_{0}^{\theta,i,N}\sim f_{0}^{\theta}$. Let $f^{\theta}_{t}$ be the unique solution to the mean-field Fokker-Planck equation (6) with initial condition $f_{0}^{\theta}=f_{0}$. Then, the $N$-particle system (4) converges to the mean-field model (5) as $N\to\infty$. That is, for any $T>0$, the $N$-particle distribution $f_{t}^{\theta,N}=\text{Law}(\mathbf{X}_{t}^{\theta,1,N},\dots,\mathbf{X}_{t}^{\theta,N,N})$ is $f_{t}$-chaotic, satisfying:
@@ 2.2 Learning Objective and Optimization
Similar to Section 2.1.2 in [68], when we only observe discrete-time samples of $M$ trajectories of an $N$-particle system at $t_{l}=l\Delta t$: $\{\mathbf{X}^{i}_{t_{l},m}\}_{i=1}^{N}$ for $m=1,\dots,M$ and $l=0,\dots,L$. Approximating the stochastic integral by the Euler–Maruyama method, we obtain the discrete-time likelihood (up to $\theta$-independent constants):
@@ 2.3 Numerical Result
We begin by validating our framework on the Motsch-Tadmor model [57, 58]. Unlike standard alignment models with additive forces, the Motsch-Tadmor dynamics feature a normalized interaction mechanism, where the influence of neighbors is weighted by their relative distance and normalized by the total interaction strength. The evolution of $N$ agents with scalar states $X^{i}_{t}\in\mathbb{R}$ is governed by:
@@
Comparisons with Gaussian Process: Additionally, we compare the learned MVNN dynamics to predictions generated from a Gaussian process model with a Matérn kernel [22]. For the comparison, the Gaussian process model was trained on $N=16$ agents and $M=9$ independent trajectories, and the system was discretized using a forward Euler scheme with $L=20$ timesteps of size $\Delta t=5\times 10^{-2}$ on the interval $[0,1]$. The MVNN model was trained on $N=16,000$ agents and $M=100$ trajectories, and the system was discretized using a forward Euler scheme with $L=200$ timesteps of size $\Delta t=10^{-2}$ on the interval $[0,2]$. The Gaussian process model is restricted to smaller training set sizes relative to the MVNN model due to the construction of a kernel matrix that scales with the total size of the training set. The initial distributions of each trajectory for both models are sampled from a Gaussian mixture model with 22-88 components, whose mean is uniformly distributed in $[0,3]$, variance is 0.25, and mixture weights are sampled from a symmetric Dirichlet distribution. Figure 2 compares the time evolution of the densities predicted by the MVNN and Gaussian process models against the ground truth for $N=16,000$ agents and three unseen initial conditions. Figure 3 shows the $L^{2}$ error between the true density and the predictions using the Gaussian process model, the MVNN model trained on 16 agents, 9 trajectories, and 20 timesteps (the same training set as the Gaussian process model), and the MVNN model trained on 16,000 agents, 100 trajectories, and 200 timesteps. The results show that MVNN matches the reference dynamics better than the Gaussian process model and achieves lower test error even when trained on a small number of agents and trajectories. MVNN is able to handle a large number of agents, which may be needed in the mean-field limit, and can also accurately account for the normalization factor in the Motsch-Tadmor model. In contrast, the Gaussian process model is limited to a smaller number of agents, and it cannot accurately handle the normalization factor, since it requires the normalized interaction kernel to be symmetric. Figure 4 compares the average simulation times against the number of agents $N$ for the MVNN and Gausian process models over 10 trials. The simulation times for MVNN remain approximately constant as $N$ increases, whereas the simulation times for the Gaussian process model increase significantly as $N$ increases.
@@
Figure 4: Simulation Time Comparison: Comparison of average simulation times (seconds) against number of agents $N$ for the MVNN and Gaussian process models over 10 trials.
@@
We further evaluate the framework on a two-dimensional first-order swarm model governed by attractive-repulsive interactions. Unlike the consensus models discussed previously, this system exhibits rich spatial pattern formation, such as rings and clumps. The dynamics of the $N$ agents are described by:
@@
where $\mathbf{X}^{i}_{t}\in\mathbb{R}^{2}$ denotes the position of agent $i$. The interaction kernel $\phi$ combines short-range repulsion and long-range attraction, modeled by a sum of Gaussians:
@@ 4 Second-Order Dynamics
where $f_{t}=\mathcal{L}\left(\mathbf{X}_{t},\mathbf{V}_{t}\right)$ describes the law of the pair $\left(\mathbf{X}_{t},\mathbf{V}_{t}\right)$, and $\mathbf{B}_{t}$ is a $d\text{-}$dimensional Wiener process. We can approximate the second-order McKean-Vlasov dynamics by a system of $N$ stochastic differential equations involving the position-velocity pairs $\left(\mathbf{X}_{t}^{i,N},\mathbf{V}_{t}^{i,N}\right)_{1\leq i\leq N}$, where $N$ is sufficiently large. The initial conditions $\left(\mathbf{X}_{0}^{i,N},\mathbf{V}_{0}^{i,N}\right)_{1\leq i\leq N}$ are independent and identically distributed with law $\mu_{0}$ and each particle evolves according to:
@@ 4.1 Measure-Valued Neural Network
The dynamics of the corresponding $N$-particle system are described by:
@@ Acknowledgments
The proof is based on a synchronous coupling argument [73]. We construct the $N$-particle system and a set of $N$ independent “ideal” mean-field processes driven by the same Brownian motions and show that their $L^{2}$ distance vanishes as $N\to\infty$. Let us define N independent processes $\bar{\mathcal{X}}^{\theta,N}_{t}=\left(\bar{\mathbf{X}}^{\theta,1,N}_{t},\cdots,\bar{\mathbf{X}}^{\theta,N,N}_{t}\right)$ defined as the solution of $N$ SDEs:
@@
where $(\mathbf{B}^{i}_{t})$ is the same Brownian motion and where we call that $f^{\theta}_{t}=\mathrm{Law}(\mathbf{X}_{t}^{\theta,i})$. Since the Brownian motions are independent, the law $f^{\theta}_{t}$ is independent of index $i$. We will show that:
@@
We partition $\Omega_{h}$ into $N^{d_{1}}$ subcubes for some $N$ to be specific later. We are going to approximate $h$ on each cube by a constnat function and then assemble them together to get an approximation of $h$ on $\Omega_{h}$. Denote the centers of the subcubes by $\{\mathbf{c}_{k}\}_{k=1}^{N^{d_{1}}}$ with $\mathbf{c}_{k}=[c_{k,1},\cdots,c_{k,d_{1}}]^{\top}$. Let $\{\mathbf{c}_{k}\}_{k=1}^{N^{d_{1}}}$ be a unifom grid on $\Omega_{h}$ so that each $\mathbf{c}_{k}\in\left\{-\gamma_{1},-\gamma_{1}+\frac{2\gamma_{1}}{N-1},\cdots,\gamma_{1}\right\}^{d_{1}}$ for each $k$. Define

### 06 Apr 2026 v2

@@ 1 Introduction
Interacting particle and agent systems are widely used to model collective dynamics in physics, biology, and the social sciences [10, 1, 35, 74, 69]. Since first-principles approaches to modeling interacting forces are often challenging, a key problem is to infer the underlying interaction mechanisms from trajectory observations. Recent data-driven methods address this by assuming a binary-interaction ansatz and estimating a pairwise interaction kernel via least-squares fitting of observed (finite-difference) velocities as a function of pairwise distances [47, 43, 37, 26, 19, 51]. However, the pairwise assumption is often insufficient for complex, realistic systems. In many biological, social, and physical scenarios, the effective dynamics are governed by a drift of the mean field form, a nonlinear functional of the population distribution, rather than a superposition of two-body forces. Popular examples include crowd dynamics governed by local density constraints rather than individual repulsion [52], vehicular traffic where speed is a nonlinear functional of the flow density [40], and cell migration influenced by chemical concentration fields [33]. Restricting the hypothesis class to pairwise interaction kernels can therefore limit the ability to capture emergent collective behaviors. Beyond modeling fidelity, computational efficiency is another major difficulty since direct simulation of interacting particle systems typically scales as $O(N^{2})$ in the number of particles. Mean-field limits and related numerical techniques, such as the random batch method [30, 23], provide scalable approximations for large populations [32, 25, 70] and have been used in aggregation and swarming, Motsch-Tadmor dynamics, and optimization [6, 13, 12, 2, 58, 11, 14, 29, 49]. This motivates the development of drift models whose evaluation scales linearly with NN, by utilizing the symmetry arising from agent indistinguishability.
@@
where $f_{t}=\mathcal{L}(\mathbf{X}_{t})$ describes the law of the $\mathbf{X}_{t}$ and $\mathbf{B}_{t}$ is a $d\text{-}$dimensional Wiener process. Numerically, the McKean–Vlasov dynamics can be approximated by a system of interacting particles $(\mathbf{X}_{t}^{i,N})_{1\leq i\leq N}$ with a sufficiently large number of particles NN. Formally, the initial condition $(\mathbf{X}_{0}^{i,N})_{1\leq i\leq N}$ are independent and identically distributed with law $\mu_{0}$ and each particle evolves according to:
@@ 2 Learning McKean–Vlasov Drifts from Particle Data
We consider a standard interacting particle/agent system as a motivating example. Note that the learning framework developed here does not rely on the pairwise interaction kernel and can be applied to more complex interactions. Consider NN agents with states $\mathbf{X}_{t}^{i}\in\mathbb{R}^{d}$ and the (possibly stochastic) dynamics:
@@ 2.1 Measure-Valued Neural Network
for some test functions $g_{i}$ and a finite-dimensional map $f$ (e.g. polynomial/smooth), see, e.g., [27, 61]. On compact subsets of Wasserstein space, such cylindrical classes are dense in appropriate smoothness spaces; see Lemma 3.12 and Definition 2.4 in [27] for a representative result. The key computational advantage is that $\langle g,\mu_{t}^{N}\rangle$ can be evaluated by averaging over particles, which is permutation invariant and scales linearly in NN. Motivated by this formulation, we approximate the drift by a composition of two neural networks: (1) an embedding network and (2) an interaction network. We define the MVNN drift as
@@
The model jointly learns (i) a finite collection of test functions $\varphi_{\mathrm{emb}}$ that extract informative features from the measure, and (ii) an interaction map $\varphi_{\mathrm{int}}$ that aggregates these features with the local state to approximate the drift. The embedding part can be interpreted as learning an optimal, data-driven test function basis on the space of measures. The embedding network $\varphi_{\mathrm{emb}}$ plays the role of the functions $g_{i}$ in the cylindrical functional representation, but unlike polynomial or finite-element bases, it is learned from data instead of fixed a priori. Given the MVNN drift (2), the associated particle system with NN agents is governed by:
@@
Although Proposition 2.1 is a standard consequence of Lipschitz continuity for McKean-Vlasov SDEs, it guarantees that our learned MVNN drift induces a well-posed nonlinear Fokker-Planck evolution at the continuum level. This establishes our method as a rigorous framework for learning continuous governing equations directly from discretized particle-level observations, without requiring any smoothing or filtering procedures, in contrast to existing approaches [55, 21]. However, to complete the theoretical justification, we must demonstrate that the learned NN-particle system (4) converges to this identified mean-field model (5) as $N\to\infty$.
@@
Let the assumptions of Proposition 2.1 hold. Let $(\mathbf{X}_{t}^{\theta,i,N})_{i=1}^{N}$ be the NN-particle system solving (4) with $f_{0}$-chaotic initial data $\mathbf{X}_{0}^{\theta,i,N}\sim f_{0}^{\theta}$. Let $f^{\theta}_{t}$ be the unique solution to the mean-field Fokker-Planck equation (6) with initial condition $f_{0}^{\theta}=f_{0}$. Then, the NN-particle system (4) converges to the mean-field model (5) as $N\to\infty$. That is, for any $T>0$, the NN-particle distribution $f_{t}^{\theta,N}=\text{Law}(\mathbf{X}_{t}^{\theta,1,N},\dots,\mathbf{X}_{t}^{\theta,N,N})$ is $f_{t}$-chaotic, satisfying:
@@ 2.2 Learning Objective and Optimization
Similar to Section 2.1.2 in [68], when we only observe discrete-time samples of $M$ trajectories of an NN-particle system at $t_{l}=l\Delta t$: $\{\mathbf{X}^{i}_{t_{l},m}\}_{i=1}^{N}$ for $m=1,\dots,M$ and $l=0,\dots,L$. Approximating the stochastic integral by the Euler–Maruyama method, we obtain the discrete-time likelihood (up to $\theta$-independent constants):
@@ 2.3 Numerical Result
We begin by validating our framework on the Motsch-Tadmor model [57, 58]. Unlike standard alignment models with additive forces, the Motsch-Tadmor dynamics feature a normalized interaction mechanism, where the influence of neighbors is weighted by their relative distance and normalized by the total interaction strength. The evolution of NN agents with scalar states $X^{i}_{t}\in\mathbb{R}$ is governed by:
@@
Comparisons with Gaussian Process: Additionally, we compare the learned MVNN dynamics to predictions generated from a Gaussian process model with a Matérn kernel [22]. For the comparison, the Gaussian process model was trained on $N=16$ agents and $M=9$ independent trajectories, and the system was discretized using a forward Euler scheme with $L=20$ timesteps of size $\Delta t=5\times 10^{-2}$ on the interval $[0,1]$. The MVNN model was trained on $N=16,000$ agents and $M=100$ trajectories, and the system was discretized using a forward Euler scheme with $L=200$ timesteps of size $\Delta t=10^{-2}$ on the interval $[0,2]$. The Gaussian process model is restricted to smaller training set sizes relative to the MVNN model due to the construction of a kernel matrix that scales with the total size of the training set. The initial distributions of each trajectory for both models are sampled from a Gaussian mixture model with 22-88 components, whose mean is uniformly distributed in $[0,3]$, variance is 0.25, and mixture weights are sampled from a symmetric Dirichlet distribution. Figure 2 compares the time evolution of the densities predicted by the MVNN and Gaussian process models against the ground truth for $N=16,000$ agents and three unseen initial conditions. Figure 3 shows the $L^{2}$ error between the true density and the predictions using the Gaussian process model, the MVNN model trained on 16 agents, 9 trajectories, and 20 timesteps (the same training set as the Gaussian process model), and the MVNN model trained on 16,000 agents, 100 trajectories, and 200 timesteps. The results show that MVNN matches the reference dynamics better than the Gaussian process model and achieves lower test error even when trained on a small number of agents and trajectories. MVNN is able to handle a large number of agents, which may be needed in the mean-field limit, and can also accurately account for the normalization factor in the Motsch-Tadmor model. In contrast, the Gaussian process model is limited to a smaller number of agents, and it cannot accurately handle the normalization factor, since it requires the normalized interaction kernel to be symmetric. Figure 4 compares the average simulation times against the number of agents NN for the MVNN and Gausian process models over 10 trials. The simulation times for MVNN remain approximately constant as NN increases, whereas the simulation times for the Gaussian process model increase significantly as NN increases.
@@
Figure 4: Simulation Time Comparison: Comparison of average simulation times (seconds) against number of agents NN for the MVNN and Gaussian process models over 10 trials.
@@
We further evaluate the framework on a two-dimensional first-order swarm model governed by attractive-repulsive interactions. Unlike the consensus models discussed previously, this system exhibits rich spatial pattern formation, such as rings and clumps. The dynamics of the NN agents are described by:
@@
where $\mathbf{X}^{i}_{t}\in\mathbb{R}^{2}$ denotes the position of agent ii. The interaction kernel $\phi$ combines short-range repulsion and long-range attraction, modeled by a sum of Gaussians:
@@ 4 Second-Order Dynamics
where $f_{t}=\mathcal{L}\left(\mathbf{X}_{t},\mathbf{V}_{t}\right)$ describes the law of the pair $\left(\mathbf{X}_{t},\mathbf{V}_{t}\right)$, and $\mathbf{B}_{t}$ is a $d\text{-}$dimensional Wiener process. We can approximate the second-order McKean-Vlasov dynamics by a system of NN stochastic differential equations involving the position-velocity pairs $\left(\mathbf{X}_{t}^{i,N},\mathbf{V}_{t}^{i,N}\right)_{1\leq i\leq N}$, where NN is sufficiently large. The initial conditions $\left(\mathbf{X}_{0}^{i,N},\mathbf{V}_{0}^{i,N}\right)_{1\leq i\leq N}$ are independent and identically distributed with law $\mu_{0}$ and each particle evolves according to:
@@ 4.1 Measure-Valued Neural Network
The dynamics of the corresponding NN-particle system are described by:
@@ Acknowledgments
The proof is based on a synchronous coupling argument [73]. We construct the NN-particle system and a set of NN independent “ideal” mean-field processes driven by the same Brownian motions and show that their $L^{2}$ distance vanishes as $N\to\infty$. Let us define N independent processes $\bar{\mathcal{X}}^{\theta,N}_{t}=\left(\bar{\mathbf{X}}^{\theta,1,N}_{t},\cdots,\bar{\mathbf{X}}^{\theta,N,N}_{t}\right)$ defined as the solution of NN SDEs:
@@
where $(\mathbf{B}^{i}_{t})$ is the same Brownian motion and where we call that $f^{\theta}_{t}=\mathrm{Law}(\mathbf{X}_{t}^{\theta,i})$. Since the Brownian motions are independent, the law $f^{\theta}_{t}$ is independent of index ii. We will show that:
@@
We partition $\Omega_{h}$ into $N^{d_{1}}$ subcubes for some NN to be specific later. We are going to approximate $h$ on each cube by a constnat function and then assemble them together to get an approximation of $h$ on $\Omega_{h}$. Denote the centers of the subcubes by $\{\mathbf{c}_{k}\}_{k=1}^{N^{d_{1}}}$ with $\mathbf{c}_{k}=[c_{k,1},\cdots,c_{k,d_{1}}]^{\top}$. Let $\{\mathbf{c}_{k}\}_{k=1}^{N^{d_{1}}}$ be a unifom grid on $\Omega_{h}$ so that each $\mathbf{c}_{k}\in\left\{-\gamma_{1},-\gamma_{1}+\frac{2\gamma_{1}}{N-1},\cdots,\gamma_{1}\right\}^{d_{1}}$ for each $k$. Define